# Towards Universal Performance Modeling for Machine Learning Training on Multi-GPU Platforms

### v2

@@ IV-B1 DLRM on Multi-GPU Platforms
Figure 10: Prediction, baseline, and reference of multi-GPU training performance of DLRMs on the two multi-GPU platforms. Notice that the “Task_x”s on different platforms do not have the same embedding table configuration as the random task generation is platform-specific. Tasks 0–9 and 10–19 on each platform are from the 2021 and 2022 parts of the dataset, respectively. Heavy tasks are marked with purple color and italics. Percentage figures are the error of bars ending in the broken area.
@@
Figure 11: Prediction, baseline, and reference of multi-GPU training performance of BERT, GPT2, and XLNet on the two multi-GPU platforms. Batch sizes greater than 32 on BERT and 16 on XLNet result in out-of-memory errors and these tests are thus skipped.
IV-B2 NLP Models Performance Modeling
@@
We also test our E2E performance model on Transformer-based NLP models including BERT, GPT2, and XLNet. In Figure 11, we see that the absolute prediction errors are less than 10% in all tests except two, which are slightly larger. The geomean prediction error of all the presented tests is 3.00%. Also, the variance of the prediction errors is obviously lower than that of DLRM workloads. The reasons are 1) these NLP models are compute (GEMM) dominated with the communication stream being well-overlapped by the compute stream, and 2) the loads across devices are more balanced than the DLRM workloads, thus intra-rank and inter-rank synchronizations are rare and have little interference in the prediction. With our highly accurate kernel performance models, we are able to precisely predict the aggregation of compute kernel time and subsequently the E2E time of these models.
V Application and Discussion
@@
V-A Case Study: Fast Sharding Config Selection Using Performance Modeling
@@
Industrial DLRM models can take days to train. Therefore, selecting a sharding configuration (i.e., the way to distribute embedding tables to multiple GPUs) that balances the loads on GPUs and speeds up the E2E per-iteration time is critical to reducing their training costs. Industrial sharding configs might consist of a sharding algorithm (greedy, multi-cost greedy, etc.), cost functions (multi-cost, memory-based, compute-based, etc.), table partition (column-based or row-based), and memory placement of tables (HBM/UVM). Previously, the best configs were selected by benchmarking and grid-searching over a big search space formed by these factors, which can take as long as one day per workload. We consider using our performance model for this task so that without running the workload, the selection time can be shortened to around 1 minute, with E2E per-iteration time predicted in seconds for each config. Notice again that we are not proposing a smart sharding algorithm here like some previous works did [27, 28, 26]. Instead, our goal is to quickly evaluate these sharding algorithms or configs and pick the fastest one for a specific problem size without benchmarking the model.
@@
We conduct an experiment to demonstrate how our performance model can quickly select the best sharding config for DLRM training on a multi-GPU system. To demonstrate the idea with simplicity, we only consider sharding algorithms as the config and omit all other factors mentioned above. We also do not include recently sophisticated sharders [27, 28, 26], although it is straightforward to integrate and test them in the future from an engineering point of view. We use six sharders listed in Table III for the experiment. In addition, we pick the 10 heavy tasks from the 20 tasks on each device generated in Section IV-B1 and run them again with the batch size 4096 and all sharders except for size_lookup_greedy. This is to guarantee that embedding lookup latency dominates the E2E time and thus the sharding config is likely to make a difference.
@@
Table IV shows the actual and prediction time of using different sharding configs in selected tasks trained with platforms 4xGV100 and 4xA100. We set the success criterion to be either the performance model accurately selects the fastest config, or the absolute error between the actual time of the predicted fastest config and the time of the actual fastest config is less than 10%. The reason for this criterion is that we not only care about if the fastest config is selected, but also how close the actual time of the predicted fastest config is from the actual fastest time when the fastest config is not selected. This is because in practice it is tolerable to fail to select the fastest config as long as the actual time of the selected config is close enough to the actual fastest time. We can see in the rightmost column that the prediction result of our performance model meets the criterion in 17 out of 20 (85%) tasks. In the rest 3 tasks, the absolute errors are only 13.76% at most. This demonstrates that our performance model is accurate enough to aid multi-GPU training optimization with low time and compute cost. Since we have demonstrated that our performance model’s prediction error is low both generally and in individual cases, we are confident that it can also perform well in unseen future cases.
V-B Discussion
The biggest advantage of our performance model is its strong adaptability to new ML models. It is essentially an execution simulator built in the granularity of kernels and ops. Supporting new ML models is straightforward, requiring only adding the missing performance models and overhead statistics of any new kernels and ops. The library of supported kernels and ops grows as the system evolves, making it even easier to support new models later. More importantly, from the performance perspective, how the workload is executed is more fundamental than how it is constructed, because the same ML model can be trained with different strategies—e.g., DP, DDP, MP, pipelining, etc.—with completely different performance. But our performance model can still handle all these cases because the easy-to-obtain execution trace (the input to our performance model) has information on both the model architecture and the execution. This means our performance model will extend well to unseen future workloads and execution paradigms.
@@ V-B Discussion
  • Our prediction pipeline only supports ML workloads in FP32 precision. However, it can be seamlessly adapted for workloads in other precision types (FP16, INT8, etc.) by preparing kernel performance models for all ops in these types.
@@
We plan to extend our current code base to support these features.
VI Conclusion and Future Work
We extend our previous work of single-GPU performance model with two critical components for more accurate single- and multiple-GPU performance modeling: support for input-data-distribution-aware performance modeling of embedding lookup as a general enhancement, and performance modeling of communication collectives and E2E training time prediction with inter-rank and intra-rank synchronization to enable multi-GPU training performance modeling of ML workloads. We demonstrate that we achieve high prediction accuracy on various types of ML workloads such as randomly generated DLRM and NLP models, and also demonstrate the performance model’s ability to speed up DLRM training through the use case of embedding table sharding config selection.
There are a series of future works to be derived from this work. Beyond the issues identified in Section V-B, the data-distribution-aware method we use to model the performance of FBGEMM embedding lookup can be generalized to other sparse ops such as SpMM and SpGEMM to handle the training and inference performance prediction of future sparsified neural networks.
@@ Acknowledgement
We would like to sincerely thank Daochen Zha, Valentin Andrei, and Yufei Zhu for their constructive feedback on this work.

### v3

@@ V-A Kernel Performance Modeling
Figure 10: Prediction, baseline, and reference of multi-GPU training performance of DLRMs on two multi-GPU platforms. Notice that the “Task_x”s on different platforms do not have the same embedding table configuration as the random task generation is platform-specific. Tasks 0–9 and 10–19 on each platform are from the 2021 and 2022 parts of the dataset, respectively. Heavy tasks are marked with purple color and italics. Percentage figures are the error of bars ending in the broken area.
V-B Multi-GPU E2E Performance Modeling
V-B1 DLRM on Multi-GPU Platforms
In our experiments to predict the multi-GPU training performance of DLRM models, we first generate 20 tasks with embedding tables randomly sampled from the DLRM open-source dataset. Among these 20 tasks, each of the 2021 and 2022 parts of the dataset contributes 5 heavy tasks (i.e., all sampled tables are heavy) and 5 normal tasks (i.e., tables can either be heavy or light). The total number of embedding tables per task is within the range of $(0.7\sim 1.3)\times\#\text{GPU}\times 13$ to resemble real-world workloads. To prevent out-of-memory errors, we limit the total memory footprint of all embedding tables on each rank to less than 80% DRAM size of the GPU for each task. We set the embedding table sharder to size_lookup_greedy based, i.e., the cost of each table is estimated as $L\times D\times\log(E)$. For each training iteration in the E2E test of each task, we randomly sample a mini-batch of data (with batch size set to 512/1024/2048/4096) corresponding to each table selected in the task from the dataset and distribute it to each rank based on the sharding scheme. The overhead statistics (mean latency, etc) of PyTorch ops are aggregated from all collected traces and shared by all tested workloads. The actual/predicted time is measured/calculated by averaging over 30 iterations. It takes about a day to run the benchmark and analysis for all 80 (20 workloads $\times$ 4 batch sizes) tasks, while it takes less than 20 minutes to predict their performance.
We present the statistics of all E2E performance modeling tests on the two multi-GPU platforms in Table II, and the prediction error and reference time of each task in Figure 10. The baseline result compared with our prediction in Figure 10 is given by the maximum sum of kernel active time of each GPU stream. We can see that this baseline prediction, yielding error values higher than 60%, is insufficient to be used as the predicted E2E time per iteration because no idle or waiting time caused by data dependency between the streams is considered. Instead, our enhanced algorithm can accurately predict both normal and heavy tasks with high accuracy, with an overall geomean prediction error of 5.21%. Most prediction results on both devices underestimate the actual time, possibly because the communication time dominates the per-iteration time when batch size is big, so the syncing and waiting time among all ranks also increases and contributes to the per-iteration time. The remaining tests that overestimate might be explained by the overestimation of GPU idle time caused by CPU overheads when the workload is latency-bound. Device-wise, the behavior of prediction errors, such as geomean/minimum/maximum error and trends when the batch size changes, does not deviate much, which justifies the consistency and stability of our prediction algorithm across platforms.
@@ V-B1 DLRM on Multi-GPU Platforms
Figure 11: Prediction, baseline, and reference of multi-GPU training performance of BERT, GPT2, and XLNet on two multi-GPU platforms. Batch sizes greater than 32 on BERT and 16 on XLNet result in out-of-memory errors and these tests are thus skipped.
V-B2 NLP Models Performance Modeling
@@
We also test our E2E performance model on Transformer-based NLP models including BERT, GPT2, and XLNet. In Figure 11, we see that the absolute prediction errors are less than 10% in all tests except two, which are slightly larger. The geomean prediction error of all the presented tests is 3.00%. Also, the variance of the prediction errors is lower than that of DLRM workloads. The reasons are 1) these NLP models are compute (GEMM) dominated with the communication stream being well-overlapped by the compute stream, and 2) the loads across devices are more balanced than the DLRM workloads, thus intra-rank and inter-rank synchronizations are rare and have little interference in the prediction. With our highly accurate kernel performance models, we can precisely predict the aggregation of compute kernel time and subsequently the E2E time of these models.
VI Application and Discussion
@@
VI-A Case Study: Fast Sharding Config Selection Using Performance Modeling
@@
Industrial DLRM models can take days to train. Therefore, selecting a sharding configuration (i.e., the way to distribute embedding tables to multiple GPUs) that balances the loads on GPUs and speeds up the E2E per-iteration time is critical to reducing their training costs. Industrial sharding configs might consist of a sharding algorithm (greedy, multi-cost greedy, etc.), cost functions (multi-cost, memory-based, compute-based, etc.), table partition (column-based or row-based), and memory placement of tables (HBM/UVM). Previously, the best configs were selected by benchmarking and grid-searching over a big search space formed by these factors, which can take as long as one day per workload. We consider using our performance model for this task so that without running the workload, the selection time can be shortened to around 1 minute, with E2E per-iteration time predicted in seconds for each config. Notice again that we are not, like some previous works did [27, 28, 26], proposing a smart sharding algorithm here to achieve optimal performance. Instead, the goal is to quickly evaluate various sharding algorithms or configs and pick the fastest one for a specific problem size without benchmarking the model.
@@
We conduct an experiment to demonstrate how our performance model can quickly select the best sharding config for DLRM training on a multi-GPU system. To show our idea with simplicity, we only consider sharding algorithms as the config and omit all other factors mentioned above. We also exclude recently sophisticated sharders [27, 28, 26], although it is straightforward to integrate and test them in the future from an engineering point of view. We use six sharders listed in Table III for the experiment. In addition, we pick the 10 heavy tasks from the 20 tasks on each device generated in Section V-B1 and run them again with the batch size 4096 and all sharders except for size_lookup_greedy. This is to guarantee that embedding lookup latency dominates the E2E time and thus the sharding config is likely to make a difference.
@@
Table IV shows the actual and prediction time of using different sharding configs in selected tasks trained with platforms 4xGV100 and 4xA100. We set the success criterion to be either the performance model accurately selects the fastest config, or the absolute error between the actual time of the predicted fastest config and the time of the actual fastest config is less than 10%. The reason for using this criterion is that we not only care about whether the fastest config is selected but also how close the actual time of the predicted fastest config is from the actual fastest time when the fastest config is not selected. This is because in practice failing to select the fastest config is tolerable as long as the actual time of the selected config is close enough to the actual fastest time. In the rightmost column, we see that the prediction result of our performance model meets the criterion in 17 out of 20 (85%) tasks. In the 3 remaining tasks, the absolute errors are only 13.76% at most. This demonstrates that our performance model is accurate enough to aid multi-GPU training optimization with low time and compute cost. Since we have demonstrated that our performance model has low prediction errors both generally and in individual cases, we are confident that it can also perform well in unseen future cases.
VI-B Discussion
The biggest advantage of our performance model is its strong adaptability to new ML models. It is essentially an execution simulator built in the granularity of kernels and ops. Supporting new ML models is straightforward, requiring only adding the missing performance models and overhead statistics of any new kernels/ops. The library of supported kernels/ops grows as the system evolves, making it even easier to support new models later. More importantly, from the performance perspective, how the workload is executed is more fundamental than how it is constructed because the same ML model can be trained with different strategies—e.g., data-parallel (DP), DDP, model-parallel (MP), pipelining, etc.—with completely different performance. However, our performance model can still handle all these cases because the easy-to-obtain execution trace (the input to our performance model) has information on both the model architecture and the execution. This means our performance model will extend well to unseen future workloads and execution paradigms.
@@ VI-B Discussion
  • Our prediction pipeline only supports ML workloads in FP32 precision. However, it can be seamlessly adapted for workloads in other precision types (FP16, INT8, etc.) by preparing kernel performance models for all kernels/ops with these types.
@@
  • The training data loading from data centers, usually impacted by uncontrollable factors like network speed and data center setups, might cause load imbalance in the industrial environment and is not considered in this work.
@@
We plan to extend our code base to support these features.
VII Conclusion and Future Work
We extend our previous work on single-GPU performance model with two critical components for more accurate single- and multiple-GPU performance modeling: performance modeling of communication collectives and E2E training time prediction with inter-rank and intra-rank synchronization to enable multi-GPU training performance modeling of ML workloads, and support for input-data-distribution-aware performance modeling of embedding lookup as a general enhancement. We achieve high prediction accuracy on various types of ML workloads such as randomly generated DLRM and NLP models, and also demonstrate the performance model’s ability to speed up DLRM training through the use case of embedding table sharding config selection.
There are a series of future works. Beyond the issues identified in Section VI-B, the data-distribution-aware method we use to model the performance of FBGEMM embedding lookup can be generalized to other sparse ops such as SpMM and SpGEMM to handle the training and inference performance prediction of future sparsified neural networks.
@@ Acknowledgement
We sincerely thank Daochen Zha, Valentin Andrei, and Yufei Zhu for their constructive feedback on this work.
Biography
[Uncaptioned image] Zhongyi Lin is Senior Software Engineer at AMD, responsible for inference optimization of Large Language Models on specialized hardware for artificial intelligence. He is experienced in building high-performance machine-learning software systems for efficient training, inference, and performance modeling of language, recommendation, and vision models on devices such as GPU and AI chips. Prior to joining AMD, he received his Ph.D. degree in Electrical and Computer Engineering from University of California, Davis.
[Uncaptioned image] Ning Sun is a Senior Performance engineer at Meta. She has over 20 years of experience in developing tools to analyze performance and detect performance anomalies in very large distributed applications. At Meta, she is working on scaling AI training stack and HHVM server on multiple generations of hardware. Prior to joining Meta, she was a lead Contributor for SPECweb and SPECjAppServer while working in the Performance team at Sun/Oracle.
[Uncaptioned image] Pallab Bhattacharya is a Distinguished Engineer at Nvidia. Prior to joining Nvidia, he helped design and deploy AI Training Cluster at Meta, moving applications from CPU-based training to GPU-based training. He has extensive experience in High-Performance Communications using Infiniband and RDMA, building, debugging, and doing performance analysis of large-scale distributed and multi-thread applications.
[Uncaptioned image] Xizhou Feng received his Ph.D. degree in Computer Science from the University of South Carolina and his J.D. degree from Marquette Law School. He is currently a software engineer at Meta Platform, Inc. His research interests include model stability and training efficiency of large AI models, scalable parallel algorithms, complex system modeling, high-performance computing, and the interactions between technology and law.
[Uncaptioned image] Louis Feng is an Engineering Manager at Meta, where he spearheads a team focused on improving fleet-wide efficiency visibility. His role involves a wide spectrum of responsibilities, including telemetry, performance analysis, and modeling. Louis’s notable contributions to the open-source community include his work on PyTorch execution trace and PARAM benchmarks, which have been adopted by the MLCommon’s Chakra working group. Before his tenure at Meta, Louis served as a staff performance engineer at Intel. There, he played key roles in the development of DreamWorks MoonRay, Pixar’s RenderMan, the Embree ray tracing engine, and the nGraph AI compiler framework. Louis received 12 patents and published papers at conferences like SIGGRAPH, NeurIPS, KDD, and ISCA. His contributions have sparked innovation across various industries.
[Uncaptioned image] John Owens is the Child Family Professor of Engineering and Entrepreneurship in the Department of Electrical and Computer Engineering at the University of California, Davis, where he leads a research group with a focus on GPU computing. He is an IEEE and AAAS Fellow and a Distinguished Member of the ACM. John earned his Ph.D. in electrical engineering in 2003 from Stanford University and his B.S. in electrical engineering and computer sciences in 1995 from the University of California, Berkeley.